### v1

@@ 3 Non-Disconnecting Arc-Kayles on structured graph families
While there are very few polynomial-time algorithms for Arc-Kayles on structured graph classes, its non-disconnecting variant seems more approachable. In this section, we expand on the results in [dailly2018criticalite], where it was shown that trees are trivial to solve. Indeed, as with $\operatorname{CSG}(\{1\})$, no strategy is required, as all possible moves on a tree will end up being played, and so the Sprague-Grundy value is either 0 or 1.
First, we will show that this result on trees can be extended to tree-like graphs: unicyclic graphs and clique trees. Recall that a graph is unicyclic if it contains exactly one cycle, and that it is a clique tree (or block graph) if every biconnected component is a clique.
Theorem 7.
Non-Disconnecting Arc-Kayles can be solved in polynomial-time on unicyclic graphs.
@@
Let $G$ be a unicyclic graph. We define cycle-moves on $G$ as legal moves that remove at least one vertex of the cycle. Cycle-moves are of two types: they can remove one vertex from the cycle that has exactly one pendant leaf (along with this leaf), or they can remove two adjacent vertices of the cycle that have no neighbor outside of the cycle. Note that playing a cycle-move leaves a tree.
The algorithm works as follows. First, assume that there is at least one cycle-move. For every cycle-move, compute the outcome of Non-Disconnecting Arc-Kayles on the remaining tree. If the outcome is $\mathcal{P}$, play this move. If no cycle-move leaves a tree with outcome $\mathcal{P}$, then, the first player loses and thus plays any move. The second player then applies the same algorithm to play their winning move. Assume now that no cycle-move exists, then the players will have to play on pendant trees until cycle-moves are open or until no move is possible. Since it is possible to know in polynomial-time if a given pendant tree will be emptied or reduced to a leaf, we can determine if a cycle-move will end up being possible. If so, compute the first player who will be able to play a cycle-move and apply the algorithm above for them. Otherwise, the graph is a disjoint union of a trees, for which the outcome can be computed in polynomial-time.
To prove the validity of the algorithm, it suffices to show that $(i)$ if no cycle-move leads to a $\mathcal{P}$-position tree, then the first player has no winning move (since, if the first player can play to such a tree, they win) and $(ii)$ it is possible to know which player will be the first to play a cycle-move.
$(i)$ Assume that there is at least one cycle-move, but all cycle-moves lead to $\mathcal{N}$-positions. The first player plays on a given pendant tree. However, this has the effect of reducing the number of vertices in any tree resulting from a cycle-move by 2. Since no strategy is possible on trees, the result is a removal of one move from such trees. Hence, all cycle-moves lead to $\mathcal{P}$-positions, and thus the second player wins by playing any cycle-move.
$(ii)$ For every pendant tree, we can compute in polynomial-time the number of moves that will either empty it or reduce it to a leaf. Thus, we can compute, for every vertex and pair of adjacent vertices of the cycle, the number of moves that will open a cycle-move on it or them. We can then compute which of the two players will benefit from those cycle-moves (by the same reasoning as in $(i)$). Hence, we can know which player will play a cycle-move first and thus win. ∎
@@
Threshold graphs are a subclass of split graphs, where the vertices can be ordered as $v_{1},\ldots,v_{n}$ such that $N(v_{1})\subseteq\ldots\subseteq N(v_{n})$, where $N(u)$ denotes the neighborhood of $u$. Alternatively, threshold graphs are graphs that can be constructed by repeatedly adding isolated or universal vertices. Note that this implies that the vertices can be partitioned into a clique and an independent set, meaning threshold graphs are split graphs. (Threshold graphs are actually exactly the intersection of split graphs and cographs.) In particular, threshold graphs can be partitioned into cliques $K_{1},\ldots,K_{m}$ and independent sets $S_{1},\ldots,S_{m}$ such that all the vertices of $S_{i}$ (resp. $K_{i}$) are twins (i.e., vertices with the exact same neighborhood), $\cup_{i=1}^{m}K_{i}$ is a clique, and the vertices of $S_{i}$ are adjacent to the vertices of $\cup_{j=1}^{i}K_{i}$. The construction order is the following: for $i$ from $n$ to 1, add the vertices of $S_{i}$ as isolated vertices, then the vertices of $K_{i}$ as universal vertices. We denote a given threshold graph by $(K\cup S,E)$, where $E$ is the edge set, $K$ is the union of the cliques, and $S$ is the union of the independent sets. If a vertex can be in both $K$ and $S$, we consider that it is in $S$ (so the independent set is maximal and the clique is minimal considering the fact that they form a partition of the vertices). An example of a threshold graph is depicted in Figure 4.
$K$$S$
@@
Recall that Non-Disconnecting Arc-Kayles is PSPACE-complete on split graphs (Theorem 2), so the following result implies that there is a complexity gap between clique-twin-free threshold graphs and split graphs:
@@
Let $T=(K\cup S,E)$ be a threshold graph with $|K|=n$ and such that $K$ is twin-free. The outcome of Non-Disconnecting Arc-Kayles on $T$ can be computed in linear time. More precisely, if $n\geq 5$, then Non-Disconnecting Arc-Kayles on $T$ is outcome-equivalent to $\operatorname{SUB}(\{1,2\})$ on a heap of size $n-1$.
@@
  • If $T$ is a threshold graph with a twin-free minimum clique of order $2$, then, as we have at least two vertices in $S$, type 2 moves are impossible. Moreover, as $K$ is twin free, and must have a universal vertex, the only possible move is of type 1. After this move, the graph is a star, and we are left with the previous case.
@@
  • If $T$ is a threshold graph with a twin-free minimum clique of order $4$, then the second player wins. If the first player plays a move of type $i$, by playing move of type $(3-i)$, the graph is left with one universal vertex in the clique and at least $3$ leaves connected to it. Therefore, no moves are available, and the second player has won.
@@
Let $i$ the largest integer such that $v_{i}$ has not been removed. By inclusion of the neighborhoods, and as $K$ is a clique and the game is non-disconnecting, $v_{i}$ is a universal vertex. Therefore, any move that does not remove $v_{i}$ cannot disconnect the graph, and is legal. Therefore, while $K$ contains at least three vertices, the two types of moves are possible.
Finally, note that, even if $|S|=|K|$ at the beginning of the game, as the optimal move for the player having a winning strategy in $\operatorname{SUB}(\{1,2\})$, is to remove the complementary (i.e. playing a move of type $(3-i)$ against a move of type $i$) to its opponent after each move, in an optimal strategy, one of the three first moves will be a move in the clique, and therefore we will reach a position with $|S|\geq|K|+2$, and thus the last vertices of the clique will not be removed. ∎

### v2

@@ 3 Non-Disconnecting Arc-Kayles on structured graph families
While there are very few polynomial-time algorithms for Arc-Kayles on structured graph classes, its non-disconnecting variant seems more approachable. In this section, we expand on the results in [dailly2018criticalite], where it was shown that trees are trivial to solve. Indeed, as with $\operatorname{CSG}(\{1\})$, no strategy is required, as all possible moves on a tree will end up being played, and so the Sprague-Grundy value is either 0 or 1. The result can be computed in linear time by maintaining a representation with levels and labeled leaves ensuring that no edge has to be considered more than once, when it is removed.
First, we will show that this result on trees can be extended to tree-like graphs: graphs with bounded feedback edge number and clique trees. Recall that the feedback edge set of a graph is a set of edges that, when removed, turn the graph into a forest; the smallest size of a feedback edge set is the feedback edge number. Also recall that a graph is a clique tree (or block graph) if every biconnected component is a clique.
See 3
To prove Theorem˜3, we present the following reduction rules. A reduction rule is called safe if it outputs a graph with the same outcome as the original graph. Moreover, since we may have to consider that a move is played, we consider the decision problem “Is the outcome of Non-Disconnecting Arc-Kayles played on $G$ oo?” for $o\in\{\mathcal{N,P}\}$, to handle changes of the current player in the reduction rules.
Reduction rule 1.
If $v\in G$ has at least three leaves attached to it, remove all of them but three.
This reduction rule can easily be applied in $\mathcal{O}(|E|)$ since each edge is considered at most twice.
Lemma 8.
Reduction rule 1 is safe
@@
Let $v$ be a vertex of $G$ and let $\ell_{1},\dots,\ell_{k}$ with $k\geq 3$ be the leaves attached to it. The first time a leaf $\ell_{i}$ is played, it must be through the edge $(\ell_{i}v)$, which would disconnect the other leaves, and at least two other leaves will remain. Thus, playing any of the leaves would not be a legal move. This statement remains true if only three leaves are attached to $v$. Thus, we can remove all but three of them without changing the legal moves, hence the outcome does not change. ∎
Let $C$ be a smallest connected subgraph of $G$ containing all the two-connected components of $G$, and let $T_{1},\dots,T_{p}$ be the connected components of $G\setminus C$. Note that each $T_{i}$ is a tree and is connected to $C$ by a single edge. Let $r_{1},\dots,r_{p}$ be the vertices of $T_{1},\dots,T_{p}$ connected to $C$.
Lemma 9.
There are at most 2k22k-2 vertices of degree at least $3$ in $C.$
Proof.
Note that $C$ has minimum degree $2$, is connected and has a feedback edge set of size $k$. Thus, we have |E|=n1+k|E|=n-1+k. Then, using the handshaking formula, we have:
$2|E|$ $=\underset{v\in C}{\sum}deg_{C}(v)$
$2(n-1+k)$ $=2n+\underset{v\in C}{\sum}(deg_{C}(v)-2)$
$2n+2k-2$ $\geq 2n+|\{v\in C|deg_{C}(v)\geq 3\}|$
$2k-2$ $\geq|\{v\in C|deg_{C}(v)\geq 3\}|$
Reduction rule 2.
Let $e_{1}=(ab),e_{2}=(cd)$ be two non-adjacent edges of $T_{i_{1}}$ and $T_{i_{2}}$ respectively (we may have $i_{1}=i_{2}$), such that the moves on $e_{1}$ and $e_{2}$ are legal (we may have that the move on $e_{2}$ is only legal after the move on $e_{1}$). We transform $G$ into $G\setminus\{a,b,c,d\}$.
$C$ can be computed in time $\mathcal{O}(|E|^{2})$ by simply starting with $C=V(G)$ and removing one by one the vertices of degree $1$ in $C$. Then this reduction rule can be applied in time $\mathcal{O}(|E|^{2})$ since each time a pair of leaves is considered, one is removed.
Lemma 10.
Reduction rule 2 is safe.
Proof.
Let $G$ be the original graph and $G^{\prime}$ be the graph where one move in $T_{i_{1}}$ and one move in $T_{i_{2}}$ have been played for some $1\leq i_{1},i_{2}\leq p$. Denote by $e_{1}=(ab)$ and $e_{2}=(cd)$ the two edges such that $G^{\prime}=G\setminus\{a,b,c,d\}$. Note that it is possible that $i_{1}=i_{2}$. We prove that $G$ and $G^{\prime}$ have the same outcome by induction on $n=|V(G)|$. If $n\leq 4$, either the graph is $K_{1,3}$ the star with three leaves, or the graph can be emptied and there is nothing to do. Suppose $n\geq 5$.
Assume first that $o(G^{\prime})=\mathcal{N}$, i.e. the first player wins in $G^{\prime}$. Let $f_{1}$ be a winning first move in $G^{\prime}$.
  • If $f_{1}\in C$, the first player also plays $f_{1}$ in $G$. Let $f_{2}$ be the answer of the second player in $G$.
    • If $f_{2}$ is a possible move in $G^{\prime}$, by the induction hypothesis, by considering $G^{\prime\prime}$ the current graph, $G^{\prime\prime}$ and $G^{\prime\prime}\setminus\{a,b,c,d\}$ have the same outcome. Thus $o(G^{\prime\prime})=\mathcal{N}$.
    • Otherwise, $f_{2}\in\{(ab),(cd)\}$. Thus, the first player can play the other edge in $\{(ab),(cd)\}$ and we are left to the game in $G$’ in which the first player has played $f_{1}$ and thus wins.
  • If $f_{1}=(a^{\prime}b^{\prime})\notin C$, then play $e_{1}$ instead. By the induction hypothesis, applied on $G\setminus\{a,b\}$ with the edges $e_{2}$ and $f_{1}$ (possible as $f_{1}$ is a legal move in $G^{\prime}$), and since $G^{\prime}\setminus\{a^{\prime},b^{\prime}\}$ has outcome $\mathcal{P}$, $G\setminus\{a,b\}$ has outcome $\mathcal{P}$, which proves that playing $e_{1}$ was a winning move for the first player.
Assume now that $o(G^{\prime})=\mathcal{P}$. Let $f_{1}=(a^{\prime}b^{\prime})$ be the first edge played in $G$.
  • If $f_{1}\in\{e_{1},e_{2}\}$, then the second player answers by playing the second edge in $\{e_{1},e_{2}\}$, and reaches the position $G^{\prime}$ which is winning for them by hypothesis.
  • If $f_{1}\notin\{e_{1},e_{2}\}$, then by the induction hypothesis applied on $G\setminus\{a^{\prime},b^{\prime}\}$, with the edges $e_{1},e_{2}$, $G\setminus\{a^{\prime},b^{\prime}\}$ has the same outcome as $G^{\prime}\setminus\{a^{\prime}b^{\prime}\}$, which is winning for the second player by hypothesis.
This proves that the reduction rule preserves the outcome and thus is safe. ∎
After applications of Reduction rule 2, at most one move can be played in a tree $T_{i}$. Up to renaming the trees, suppose it is in $T_{1}$. Note that we cannot get rid of the last edge on which a move is possible since it may change the strategy in the rest of the graph, and it might be that the outcome doesn’t change by removing an edge in a tree (see Figure 4).
uuvv
Figure 4: Example of graph having a pendant edge $(uv)$ such that $o(G)=\mathcal{N}$ and $o(G\setminus\{u,v\})=\mathcal{N}$.
Reduction rule 3.
If $F$ is a forest containing at least $2$ vertices in which no move can be played, and attached to some vertex $u$, we replace $T_{u}$ the subtree induced by $u$ and this forest by a tree $T^{\prime}$ of Figure 6 such that $o(T_{u})=o(T^{\prime})$ and $o(T_{u}-\{u\})=o(T^{\prime}-\{u\})$ (or both $T_{u}-\{u\}$ and $T^{\prime}-\{u\}$ are disconnected).
If $F$ is a forest containing at least $4$ vertices in which exactly one legal move $e$ can be played, and attached to some vertex $u$, we proceed the same way considering $F-\{e\}$ and we then add a path of length 2 attached to a leaf of $T^{\prime}$.
An example of application of Reduction rule 3 is provided in Figure 5
[…]
Figure 5: An example of application of Reduction rule 3. Here, we have $o(T_{u})=\mathcal{P}$ and $o(T_{u}-\{u\})=\mathcal{N}$.
This reduction rule can be applied in time $\mathcal{O}(|V|)$ since each vertex is in at most one forest and, as mentioned before, the outcome of the game played on a tree can be computed in linear time. If no legal move remains in $F$, we present in Figure 6 a set of trees covering all the possible outcomes.
Lemma 11.
Reduction rule 3 is safe.
In order to prove Lemma 11, we prove the following stronger lemma (unlike in Reduction rule 3, we take the root into account in the number of vertices):
Lemma 12.
Let $G_{0}$ be a graph, and let $(T,r)$ and $(T^{\prime},r^{\prime})$ be two rooted trees of order at least 33 (55 resp.), in which no (exactly one resp.) legal move not containing the root can be made. Let $u\in G$ be a vertex. Suppose $o(T)=o(T^{\prime})$, $T-\{r\}$ is disconnected if and only if $T^{\prime}-\{r^{\prime}\}$ is and $o(T-\{r\})=o(T^{\prime}-\{r^{\prime}\})$ otherwise. Let $G$ and $G^{\prime}$ be the graphs in which $T$ and $T^{\prime}$ are attached to $G_{0}$ by identifying $r$ and $u$ or $r^{\prime}$ and $u$ respectively. Then $o(G)=o(G^{\prime})$.
Proof.
Suppose that the first (second resp.) player has a winning strategy in $G$. We consider the following strategy in $G^{\prime}$. While there is a vertex in $G_{0}$, it is not possible to play any edge containing $u$ without disconnecting the remaining vertices of $G_{0}$ and the remaining vertices of $T$ or $T^{\prime}$. Thus, there is a natural bijection between the moves of $G$ and the moves of $G^{\prime}$, since there is at most one move outside $G_{0}$ that can be made in the bijection if it exists. When the last move is $G_{0}$ is made, there are only four possibilities:
  • At least three vertices remain in $G_{0}$. In this case, the game ends as it is not possible to claim $u$ and a vertex of $T$ nor $T^{\prime}$, since it would disconnect the remaining vertices of $G_{0}$ with the remaining vertices of $T$ or $T^{\prime}$, which exists since we supposed that they have order at least $5$.
  • Exactly two vertices remain in $G_{0}$. Then, since all the moves have been legal, they are exactly $u$ and a leaf $\ell$ connected to it. Since the last move of $G_{0}$ has been played, playing $(u\ell)$ is not legal, which means that the resulting graph would be $T-\{r\}$, which would be disconnected. But, by the hypothesis this means that $T^{\prime}-\{r^{\prime}\}$ is also disconnected and the game is also finished in $G^{\prime}$.
  • Exactly one vertex remains in $G_{0}$. It must be $u$ since, otherwise, $G$ would be disconnected. The game is then equivalent to the game played on $T$, which has the same outcome as the game played on $T^{\prime}$
  • No vertex remains in $G_{0}$. The game is then equivalent to the game played on $T-\{r\}$, which has the same outcome as the game played on $T^{\prime}-\{r^{\prime}\}$. ∎
$u$
(a) A tree $T$ satisfying $o(T)=\mathcal{N}$ where removing $u$ disconnects the graph.
$u$
(b) A tree $T$ satisfying $o(T)=\mathcal{P}$ where removing $u$ disconnects the graph.
$u$
(c) A tree $T$ satisfying $o(T)=\mathcal{N}$ and
$o(T-\{u\})=\mathcal{N}$.
$u$
(d) A tree $T$ satisfying $o(T)=\mathcal{P}$ and
$o(T-\{u\})=\mathcal{N}$.
$u$
(e) A tree $T$ satisfying $o(T)=\mathcal{P}$ and
$o(T-\{u\})=\mathcal{P}$.
$u$
(f) A tree $T$ satisfying $o(T)=\mathcal{N}$ and
$o(T-\{u\})=\mathcal{P}$.
Figure 6: A set of small trees with different outcomes.
After an exhaustive use of Reduction rule 3, the graph consists in a core $C$ of vertices of degree at least 2 and containing all the 2-connected components of the graph and trees of size at most $10$ attached to it. Recall that, using Lemma 9, there are at most 2k22k-2 vertices having at least $3$ neighbors in $C$. Thus, we only have to reduce long paths of vertices having $2$ neighbors in $C$. To achieve this, we give types to the vertices of $C$ as follows:
  • type 0: if nothing is attached to it
  • type 1: if a leaf is attached to it.
  • type M: if a tree in which there is a legal move from a leaf attached to it.
  • type B1: if exactly three leaves are attached to it.
  • type B2: if any other tree is attached to it.
We also denote by $M$ the set of vertices of type M, and by $B$ the set of vertices of type B1 or B2. Note that, after exhausting Reduction rule 2, there is at most one vertex of type M. Let $S=\{v\in C|deg_{C}(v)\geq 3\}\cup M$. Using Lemma 9, we have that $|S|\leq 2k-1$.
We now prove that there exists a constant $\alpha$ such that we can reduce any path of vertices of type 0, 1 or $B$ to a path of length at most $\alpha$
Let $P$ be a path of vertices of type 0, 1 or B, and of length at least $3$. Let $a,b$ be the two vertices of $S$ connected to the endpoints of $P$, denoted by $u$ and $v$ respectively. For the rest of the proof, we also consider a new outcome $\mathcal{X}$ corresponding to “the move is illegal”. Let $\mathcal{O}=\{\mathcal{N},\mathcal{P}\}$ the classical set of outcome and $\mathcal{O}^{\prime}=\mathcal{O}\cup\{\mathcal{X}\}$. We define the extended outcome $\tilde{o}(P)=(o_{1}\times\varepsilon_{1},o_{2}\times\varepsilon_{2},o_{3}\times\varepsilon_{3},o_{4}\times\varepsilon_{4},o_{5}\times\varepsilon_{5},O\times\varepsilon_{6})$, with $o_{1},o_{2},o_{3},o_{4},o_{5}\in\mathcal{O}^{\prime}$ the outcome of the subtree where $a,b,\{a,u\},\{b,v\},\{a,b\}$ are removed respectively, $O\subset 2^{\mathcal{O}\times\mathcal{O}}$ the set of pairs of outcomes possible after a legal move in $P$ on the subtree rooted in $v$ and the one rooted in $u$, and $\varepsilon_{1}\dots,\varepsilon_{6}\in\{0,1\}$ the parity of the number in the remaining trees. Note that $\tilde{o}(P)$ can only take at most $3^{5}*2^{4}*2^{6}=248832$ different values.
Let $\mathcal{P}$ be a finite set of paths (with a forest attached to them) such that $\tilde{o}(\mathcal{P})$ covers all the possible outcomes of $\tilde{o}$.
Note that, in practice, these paths are paths in $C$, but are trees in $G$.
Reduction rule 4.
While there is a path of vertices of type $0,1$ or B of length at least $3$, that is not in $\mathcal{P}$ in $G$, replace it with one of $\mathcal{P}$ with the same extended outcome.
This reduction rule can be computed in time $\mathcal{O}(|E|^{2})$ as for each of the graphs, the outcome of two trees has to be computed. The following Lemma can be proved using the same idea as Lemma 11, but with more case analysis.
Lemma 13.
Reduction rule 4 is safe.
Proof.
Let $G^{\prime}$ be the graph obtained from $G$ by replacing a path $P$ by a path $P^{\prime}$ of $\mathcal{P}$ having the same extended outcome. Let $a,b$ be the two vertices of $C$ that are connected to $P$. Let $u,v$ be the vertices of $P$ connected to $a$ and $b$ respectively in $G$ and $u^{\prime},v^{\prime}$ be the vertices of $P^{\prime}$ connected to $a$ and $b$ respectively in $G^{\prime}$
Suppose that the first (second resp.) player has a winning strategy in $G$, we consider the following strategy in $G^{\prime}$. While the optimal move in $G$ is in $G\setminus(P\cup\{a,b\})$ and while the opponents answers in $G^{\prime}\setminus(P^{\prime}\cup\{a,b\})$ , play the same edge in $G^{\prime}$ or consider he answers the same edge in $G\setminus(P\cup\{a,b\})$. Consider the first time another move is played:
  • If the opponent plays an edge in $G^{\prime}\setminus P^{\prime}$ containing $a$ or $b$, consider that he played the same move in $G\setminus P$
  • If the opponent plays the $(au^{\prime})$ or $(bv^{\prime})$, consider that he played , $(au)$ or $(bv)$ respectively.
  • If the opponent plays an edge in $P^{\prime}$, let $o,o^{\prime}$ be the outcome of the subtrees rooted in $u^{\prime}$ and $v^{\prime}$, and consider that he plays an edge in $P$ such that the subtrees rooted in $u$ and $v$ in $G$ have the same outcome
  • If the optimal move is to play an edge in $G\setminus P$ containing $a$ or $b$, play the same move in $G^{\prime}\setminus P^{\prime}$
  • If the optimal move is to play $(au)$ or $(bv)$, play $(au^{\prime})$ or $(bv^{\prime})$ respectively.
  • If the optimal move is to play an edge in $P$, let $o,o^{\prime}$ be the outcome of the subtrees rooted in $u$ and $v$, and play an edge in $P^{\prime}$ such that the subtrees rooted in $u^{\prime}$ and $v^{\prime}$ in $G^{\prime}$ have the same outcome.
After these moves, all the remaining vertices of $P$ and $P^{\prime}$ are trees attached to vertices of $C$. Thus, using Reduction rule 2, we can remove moves in them such that at most one move is playable in each of them (and one remains in $P$ if and only if one remains in $P^{\prime}$). Thus, we can now apply Reduction rule 3 to replace both of them by the same tree, which results into the same graph. Overall, the resulting position if equivalent to the same graph, say $G^{\prime\prime}$. Thus, as the strategy was a winning strategy for the first (second resp.) played in $G$, it is also one in $G^{\prime}$
Proof of Theorem˜3.
Let $\alpha$ be the maximum size of the trees in $\mathcal{P}$. By exhausting application of Reduction rule 4, the resulting graph has at most (2k1)(2k-1) vertices of degree at least $3$ in $C$, all of them are connected through paths of lengths at most $\alpha$, and each of these vertices is attached to trees of order at most $10$. Overall, after applications of these rules, there are at most $40\alpha k^{2}+\mathcal{O}(k)$ vertices in the resulting graph which has the same outcome as $G$. Thus, it is a quadratic kernel. Moreover, since all the the reduction rules can be computed in quadratic time, we can compute this kernel in quadratic time and compute the winner in time $\mathcal{O}(|E|^{2}f(k))$. ∎
We now study more structured graph classes to obtain efficient algorithms: we provide polynomial-time algorithms for clique trees and a subclass of threshold graphs.
A clique tree is a tree where every vertex is expanded into a clique (which can be of order 1).
@@
Threshold graphs are a subclass of split graphs, where the vertices can be ordered as $v_{1},\ldots,v_{n}$ such that $N(v_{1})\subseteq\ldots\subseteq N(v_{n})$, where $N(u)$ denotes the neighborhood of $u$. Alternatively, threshold graphs are graphs that can be constructed by repeatedly adding isolated or universal vertices. Note that this implies that the vertices can be partitioned into a clique and an independent set, meaning threshold graphs are split graphs. (Threshold graphs are actually exactly the intersection of split graphs and cographs, i.e., graphs that contain no induced $P_{4}$.) In particular, threshold graphs can be partitioned into cliques $K_{1},\ldots,K_{m}$ and independent sets $S_{1},\ldots,S_{m}$ such that all the vertices of $S_{i}$ (resp. $K_{i}$) are twins (i.e., vertices with the exact same neighborhood), $\cup_{i=1}^{m}K_{i}$ is a clique, and the vertices of $S_{i}$ are adjacent to the vertices of $\cup_{j=1}^{i}K_{i}$. The construction order is the following: for $i$ from $n$ to 1, add the vertices of $S_{i}$ as isolated vertices, then the vertices of $K_{i}$ as universal vertices. We denote a given threshold graph by $(K\cup S,E)$, where $E$ is the edge set, $K$ is the union of the cliques, and $S$ is the union of the independent sets. If a vertex can be in both $K$ and $S$, we consider that it is in $S$ (so the independent set is maximal and the clique is minimal considering the fact that they form a partition of the vertices). An example of a threshold graph is depicted in Figure˜7.
KKSS
@@
Another characterization of threshold graphs is as the intersection of split graphs and cographs. Recall that Non-Disconnecting Arc-Kayles is PSPACE-complete on split graphs (Theorem˜2), so the following result implies that there is a complexity gap between clique-twin-free threshold graphs and split graphs:
@@
Let $T=(K\cup S,E)$ be a threshold graph with $|K|=n$ and such that $K$ is twin-free. If $n\geq 5$, then Non-Disconnecting Arc-Kayles on $T$ is outcome-equivalent to $\operatorname{SUB}(\{1,2\})$ on a heap of size n1n-1, and thus its outcome can be computed in polynomial time.
@@
  • If $T$ is a threshold graph with a twin-free minimum clique of order $2$, then, as we have at least two vertices in $S$, type 2 moves are impossible. Moreover, as $K$ is twin-free, and must have a universal vertex, the only possible move is of type 1. After this move, the graph is a star, and we are left with the previous case.
@@
  • If $T$ is a threshold graph with a twin-free minimum clique of order $4$, then the second player wins. If the first player plays a move of type $i$, by playing move of type (3i)(3-i), the graph is left with one universal vertex in the clique and at least $3$ leaves connected to it. Therefore, no moves are available, and the second player has won.
@@
Let $i$ be the largest integer such that $v_{i}$ has not been removed. By inclusion of the neighborhoods, and as $K$ is a clique and the game is non-disconnecting, $v_{i}$ is a universal vertex. Therefore, any move that does not remove $v_{i}$ cannot disconnect the graph, and is legal. Therefore, while $K$ contains at least three vertices, the two types of moves are possible.
Finally, note that, even if $|S|=|K|$ at the beginning of the game, as the optimal move for the player having a winning strategy in $\operatorname{SUB}(\{1,2\})$, is to remove the complementary (i.e. playing a move of type (3i)(3-i) against a move of type $i$) to its opponent after each move, in an optimal strategy, one of the three first moves will be a move in the clique, and therefore we will reach a position with $|S|\geq|K|+2$, and thus the last vertices of the clique will not be removed. ∎